OpenStudy (anonymous):

The ratio of two supplementary angles is 4:5, find the measures of each angle ?

OpenStudy (alexandervonhumboldt2):

suplmentary angles are who add to 180 degress

OpenStudy (andrewhaze):

so 2 angles =4.5

OpenStudy (alexandervonhumboldt2):

a/b=4/5 a+b=180

OpenStudy (alexandervonhumboldt2):

a=4/5b a+b=180

OpenStudy (alexandervonhumboldt2):

9/5b=180

OpenStudy (alexandervonhumboldt2):

b=100 a=80

OpenStudy (alexandervonhumboldt2):

80/100=4:5

OpenStudy (alexandervonhumboldt2):

100, 80

OpenStudy (anonymous):

Can you explain that step by step ?

OpenStudy (alexandervonhumboldt2):

suplmentary angles are who add to 180 degress create system of equations: a/b=4/5 and a+b=180. a=4/5b a+b=180. substitute a: 9/5b=180 b=100. a=180-100=80. b=100 a=80